# Fixed-domain curve counts for blow-ups of projective space

### v3

@@ 2. Case 2
    If $f_{*}[C_{{\text{sp}}}]=d^{\prime}\mathsf{H}^{\vee}+k_{1}^{\prime}\mathsf{E}_{1}^{\vee}+\cdots+k_{\ell}^{\prime}\mathsf{E}_{\ell}^{\vee}$ with $d^{\prime}>0$ or it is the 00 class, then condition (*’) follows from Lemma 22.
@@ 3.1. Integral formula on $\mathbb{P}$
In this section, we prove Proposition 10. We first recall the setup of §1.3.2.
@@
Let $S:=\mathsf{Jac^{d}}(C)\times\mathsf{Sym}^{k_{1}}(C)\times\ldots\times\mathsf{Sym}^{k_{\ell}}(C)$, and let $\nu:\mathbb{P}\to S$ be the projective bundle parametrizing $\ell$-tuples of sections
@@
We wish to cut out subschemes of $V(x_{i})\subset\mathbb{P}$ corresponding to the conditions $f(p_{i})=x_{i}$, but care must be taken when the $p_{i}$ vanishes at some of the $f_{j}$. For a given point $[f_{1}:\cdots:f_{r+1}]$ of $\mathbb{P}$, in order for $f(p_{i})=x_{i}$, we need
@@
cut out by the equations (12) for all $h,j$ with $11\leq h<j\leq a+$ 1, where we interpret $f_{h}$ and $f_{j}$ as sections of $\mathcal{L}^{\prime}(D_{h}+D_{j})$, and then restrict to $p_{i}$. That is, $V(x_{i})_{a}$ is the locus of $f$ where the first $a+1$ coordinates of $f(p_{i})$ agree with those of $x_{i}$, where the $h$-th and $j$-th coordinates of $f(p_{i})$ and $x_{i}$ are compared by regarding $f_{h},f_{j}$ as sections of $\mathcal{L}^{\prime}(D_{h}+D_{j})$.
@@
Theorem 10 is implied by the following two statements.
Proposition 33.
@@
$[V(x_{i})_{a}]=\widetilde{\mathsf{H}}^{a}+\sigma_{1}\widetilde{\mathsf{H}}^{a-1}+\cdots+\sigma_{a}\text{ in }H^{2a}(\mathbb{P}).$
@@
Proposition 34 (Transversality).
The scheme $V(x_{1},\ldots,x_{n})$ is reduced and 00-dimensional. Moreover, its $\mathbb{C}$-points are in bijection with the set of maps enumerated in $\mathsf{Tev}^{X}_{g,n,\beta}$. (In particular, the $f_{j}$, when regarded as sections of $\mathcal{L}$, do not simultaneously vanish at any $p\in C$.)
The proof of transversality is deferred to the next section. Here, we prove Proposition 33. Throughout the rest of this section, we work with a fixed $i$, and assume that $x_{i}=[1:\cdots:1]$. Also, by setting $\eta_{i}=0$ for $i>\ell$, we assume that $\ell=r+1$.
Lemma 35.
@@
is surjective, because $H^{1}(C\times B,\mathcal{L}^{\prime}(D_{1})(-(C\times 0)-(p_{i}\times B)))=0$. Note here that in order to ensure that $R^{1}\pi_{*}(\mathcal{L}^{\prime}(D_{1})(-(C\times 0)-(p_{i}\times B)))=0$, we need the right hand side of (3) to be $12g-$ 1, rather than $22g-$ 2. Therefore, we can define all other $\widetilde{f}_{t}$ extending $f_{t}$ and with $\widetilde{f}_{1}(p_{i})=\widetilde{f}_{t}(p_{i})$ in $\mathcal{L}$. $\Diamond$
@@
Lemma 37.
Let $[V]_{a}\in H^{2a}(\mathbb{P})$ be the class of $V(x_{i})_{a}\subset\mathbb{P}$. Then, for all $a\leq r$, we have
(13) $[V]_{a}=[V]_{a-1}(\widetilde{\mathsf{H}}+\eta_{a}+\eta_{a+1})-[V]_{a-2}(\widetilde{\mathsf{H}}+\eta_{a})\eta_{a},$
where we set $[V]_{-1}=0$ by convention.
@@
The claim follows from the following three statements, after applying Lemma 36.
@@
(2) is proven exactly as in Lemma 35, namely, $V_{a-1}\cap W_{a,a+1}$ is generically a sub-projective bundle of $\mathbb{P}$ of codimension $a$, and $V_{a-2}\cap W_{a}\cap N_{a}$ is generically a sub-projective bundle of the pullback of $\mathbb{P}$ over $N_{a}$ of codimension $a-1$. The details are omitted.
@@
Proof of Proposition 33.
We proceed by induction on $a$. When $a=1$, this is Lemma 36. Suppose $a>1$. By the inductive hypothesis and Lemma 37, we have
@@
$,\displaystyle(\eta_{a}+\eta_{a+1})\sigma_{i-1}(\eta_{1},\ldots,\eta_{a})+\sigma_{i}(\eta_{1},$$\ldots,\eta_{a})-\eta_{a}\sigma_{i-1}(\eta_{1},\ldots,\eta_{a-1})$
@@ 3.1.1. Interlude on the permutohedral variety
Fix a dimension $r\geq 2$. Write $[r+1]=\{1,2,\ldots,r+1\}$. For a non-empty subset $S\subsetneq[r+1]$, let $\Lambda_{S}\subset\mathbb{P}^{r}$ denote the torus-invariant linear space of dimension (#$(\#S-1)$ given by the vanishing of the coordinates indexed by the complement $[r+1]\backslash S$.
Definition 38.
@@
Definition 39.
For $S\subset[r+1]$ of cardinality at most $r-1$, let $\mathsf{E}_{S}$ be the class of the strict transform of $\Lambda_{S}$. Let $\mathsf{H}^{\vee},\mathsf{E}_{S}^{\vee}\in H_{2}(Y)$ be the basis of 1-cycles dual to the basis $\mathsf{H},\mathsf{E}_{S}\in H^{2}(Y)$ of divisors.
@@
A point $y\in Y$ can be expressed in coordinates as follows. First, take a point $y_{0}\in\mathbb{P}^{r}=\mathbb{P}(\mathbb{C}[[r+1]]),$ where $\mathbb{C}[[r+1]]$ denotes the vector space with basis given by the set $[r+1]=\{1,2,\ldots,r+1\}$. The point $y_{0}$ is the image of $y$ under the composite blow-up $Y\to\mathbb{P}^{r}$. Then, let $S^{1}\subsetneq[r+1]=:S^{0}$ be the subset of coordinates of $y_{0}$ which are equal to zero, corresponding to the minimal $T$-invariant subvariety of $\mathbb{P}^{r}$ in which $y_{0}$ lies. Then, let $y_{1}\in\mathbb{P}(\mathbb{C}[S^{1}])\cong\mathbb{P}^{\#S^{1}-1}$ be a point, representing a projectivized normal vector to $\Lambda_{(S^{1})^{c}}$ at $y_{0}$. Define $S^{2}\subsetneq S^{1}$ analogously, as the set of coordinates of $y_{1}$ equal to zero, and continue until $y_{k}$ has all coordinates non-zero. Then, $y$ consists of the data of the points $(y_{0},y_{1},\ldots,y_{k})$.
@@
Let $C$ be a smooth curve. A map $f:C\to Y$, which, upon post-composition with the blow-up $Y\to\mathbb{P}^{r}$, has image not contained in any torus-invariant subvariety, may be is given by the following data:
@@
  • for each $S\subset[r+1]$ with $11\leq\#S\leq r-$ 1, an effective divisor $D_{S}\subset C$, and
@@
The point $f(p)=(y_{0},\ldots,y_{k})$ may be computed as follows. First $y_{0}$ is the point $[f_{0}(p):\cdots:f_{r+1}(p)],$ where the $f_{j}$ are regarded as sections of $\mathcal{L}$. Then, the subset $S^{1}\subset[r+1]$ is the set of indices $j$ for which $f_{j}(p)=0$. The point $y_{1}$ has coordinates given by the values of $f_{j}(p)$ after twisting $\mathcal{L}$ down by the unique positive multiple of $p$ for which the $f_{j}(p)$ are well-defined and not all zero for $j\in S^{1}$. The rest of the $y_{2},\ldots,y_{k}$ are then determined by further twists of $\mathcal{L}$.
@@ 3.1.2. Transversality
In this section, we prove Proposition 34. We continue to assume $\ell=r+1$. The main difficulty is to show that a point of $V(x_{1},\ldots,x_{n})$ represents an “honest” map in class $\beta$. More precisely, let $\mathbb{P}^{\circ}\subset\mathbb{P}$ be the open locus of $f=[f_{1}:\cdots:f_{r+1}]$ for which the $f_{j}$ share no common zeroes, when simultaneously viewed as sections of $\mathcal{L}$. Then, we wish to show that $V(x_{1},\ldots,x_{n})\subset\mathbb{P}^{\circ}$.
@@
Remark 41.
We will later use the assumption $n-d\geq g+1$ again (see Lemma 50), but we point out that it is crucial for Lemma 40 to hold, which in turn is necessary for the transversality.
@@
On the other hand, when $g=0$ and $r=2$, one can check that the assumption $n-d\geq 1$ can be dropped, also later in Lemma 50.
We now turn to the main argument. The strategy is as follows: for any initial $f_{\text{init}}\in V(x_{1},\ldots,x_{n})$, we describe an algorithm that essentially amounts to twisting down at base-points of $f_{\text{init}}$ until we are able to define a map $f:C\to\mathbb{P}^{r}$. The map $f$ will also have various incidence conditions with respect to the torus-invariant loci of $\mathbb{P}^{r}$ and the $p_{i},x_{i}$; a naive dimension count would immediately predict the nonexistence of such a map, unless $f\in\mathbb{P}^{\circ}$ to begin with. We show that this expectation holds by passing from $f$ to a map from $f:C\to Y$, where $Y$ is the permutohedral variety of the previous section, moving in a family of the expected dimension by Lemma 1.
Our procedure consists of a sequence of modifications to the following data:
@@
  • For all $j=1,2,\ldots,r+1$, a non-zero (see Lemma 40) section
@@
The initial data is given by the data underlying the point $f_{\text{init}}\in\mathbb{P}$: namely, the underlying line bundle denoted $\mathcal{L}_{\text{init}}$, the divisors $D_{\{j\}}:=D_{j}$ and $D_{S}=0$ if #$\#S>1$, and the sections $(f_{j})_{\text{init}}$ of $\mathcal{L}_{\text{init}}(-\sum_{v\neq j}D_{j})$. We say that $p$ is a base-point for the above data if $f_{j}(p)=0$ as a section of $\mathcal{L}$ for all $j=1,2,\ldots,r+1$.
Throughout, we distinguish the initial line bundle and sections $\mathcal{L}_{\text{init}},(f_{j})_{\text{init}}$, which do not change, from the $\mathcal{L},f_{j}$, which do. The letter $f$ is only used at the end when we obtain a map $f:C\to\mathbb{P}^{r}$ after no base-points remain. Similarly, we use $D_{j}$ to denote the divisor associated to the point $f_{\text{init}}\in V(x_{1},\ldots,x_{n})$, which does not change, to distinguish it from $D_{\{j\}}$, which does.
We will repeatedly consider the following conditions on $f_{j}$ given $p\in C$:
@@ 1. Step 1
    Let $\alpha>0$ be the largest order to which $p$ vanishes at every $f_{j}$ (as a section of $\mathcal{L}(-\sum_{v\neq j}D_{\{v\}})$) to order $\alpha$. Then, make the following modifications:
@@
    • each $D_{S}$ stays the same (in particular, all $D_{S}$ with #$\#S>1$ remain empty), and
@@
      (In other words, $f_{j}$ is “twisted down by $\alpha$ at $p$.”)

### v4

@@ 2. Case 2
    If $f_{*}[C_{{\text{sp}}}]=d^{\prime}\mathsf{H}^{\vee}+k_{1}^{\prime}\mathsf{E}_{1}^{\vee}+\cdots+k_{\ell}^{\prime}\mathsf{E}_{\ell}^{\vee}$ with $d^{\prime}>0$ or it is the 0 class, then condition (*’) for sufficiently large $n$ follows from Lemma 2.7.
@@ 3.1. Integral formula on $\mathbb{P}$
In this section, we prove Proposition 1.10. We first recall the setup of §1.3.2.
@@
Let $S:=\mathsf{Jac^{d}}(C)\times\mathsf{Sym}^{k_{1}}(C)\times\ldots\times\mathsf{Sym}^{k_{\ell}}(C)$, and let $\nu:\mathbb{P}\to S$ be the projective bundle whose fiber over $(\mathcal{L},D_{1},\ldots,D_{\ell})\in S$ parametrizes $(r+1)$-tuples of sections
@@
We wish to cut out subschemes of $V(x_{i})\subset\mathbb{P}$ corresponding to the conditions $f(p_{i})=x_{i}$, but care must be taken when the $p_{i}$ vanishes at some of the $f_{j}$. For a given point $f_{1}:\cdots:f_{r+1}]$ of $\mathbb{P}$, in order for $f(p_{i})=x_{i}$, we need
@@
cut out by the equations (16) for all $h,j$ with $1\leq h<j\leq a+1$, where we interpret $f_{h}$ and $f_{j}$ as sections of $\mathcal{L}^{\prime}(D_{h}+D_{j})$, and then restrict to $p_{i}$. That is, $V(x_{i})_{a}$ is the locus of $f$ where the first $a+1$ coordinates of $f(p_{i})$ agree with those of $x_{i}$, where the $h$-th and $j$-th coordinates of $f(p_{i})$ and $x_{i}$ are compared by regarding $f_{h},f_{j}$ as sections of $\mathcal{L}^{\prime}(D_{h}+D_{j})$.
@@
Theorem 1.10 is implied by the following two statements.
Proposition 3.1.
@@
$V(x_{i})_{a}]=\widetilde{\mathsf{H}}^{a}+\sigma_{1}\widetilde{\mathsf{H}}^{a-1}+\cdots+\sigma_{a}\text{ in }H^{2a}(\mathbb{P}).$
@@
Proposition 3.2 (Transversality).
The scheme $V(x_{1},\ldots,x_{n})$ is reduced and 0-dimensional. Moreover, its $\mathbb{C}$-points are in bijection with the set of maps enumerated in $\mathsf{Tev}^{X}_{g,n,\beta}$. (In particular, the $f_{j}$, when regarded as sections of $\mathcal{L}$, do not simultaneously vanish at any $p\in C$.)
The proof of transversality is deferred to the next section. Here, we prove Proposition 3.1. Throughout the rest of this section, we work with a fixed $i$, and assume that $x_{i}=[1:\cdots:1]$. Also, by setting $\eta_{i}=0$ for $i>\ell$, we assume that $\ell=r+1$.
Lemma 3.3.
@@
is surjective, because $H^{1}(C\times B,\mathcal{L}^{\prime}(D_{1})(-(C\times 0)-(p_{i}\times B)))=0$. Note here that in order to ensure that $R^{1}\pi_{*}(\mathcal{L}^{\prime}(D_{1})(-(C\times 0)-(p_{i}\times B)))=0$, we need the right hand side of (3) to be 2g12g-1, rather than 2g22g-2. Therefore, we can define all other $\widetilde{f}_{t}$ extending $f_{t}$ and with $\widetilde{f}_{1}(p_{i})=\widetilde{f}_{t}(p_{i})$ in $\mathcal{L}$. $\Diamond$
@@
Lemma 3.5.
Let $V]_{a}\in H^{2a}(\mathbb{P})$ be the class of $V(x_{i})_{a}\subset\mathbb{P}$. Then, for all $a\leq r$, we have
(17) $V]_{a}=[V]_{a-1}(\widetilde{\mathsf{H}}+\eta_{a}+\eta_{a+1})-[V]_{a-2}(\widetilde{\mathsf{H}}+\eta_{a})\eta_{a},$
where we set $V]_{-1}=0$ by convention.
@@
The claim follows from the following three statements, after applying Lemma 3.4.
@@
(2) is proven exactly as in Lemma 3.3, namely, $V_{a-1}\cap W_{a,a+1}$ is generically a sub-projective bundle of $\mathbb{P}$ of codimension $a$, and $V_{a-2}\cap W_{a}\cap N_{a}$ is generically a sub-projective bundle of the pullback of $\mathbb{P}$ over $N_{a}$ of codimension a1a-1. The details are omitted.
@@
Proof of Proposition 3.1.
We proceed by induction on $a$. When $a=1$, this is Lemma 3.4. Suppose $a>1$. By the inductive hypothesis and Lemma 3.5, we have
@@
$(\eta_{a}+\eta_{a+1})\sigma_{i-1}(\eta_{1},\ldots,\eta_{a})+\sigma_{i}(\eta_{1},$ $\ldots,\eta_{a})-\eta_{a}\sigma_{i-1}(\eta_{1},\ldots,\eta_{a-1})$
@@ 3.1.1. Interlude on the permutohedral variety
Fix a dimension $r\geq 2$. Write $r+1]=\{1,2,\ldots,r+1\}$. For a non-empty subset $S\subsetneq[r+1]$, let $\Lambda_{S}\subset\mathbb{P}^{r}$ denote the torus-invariant linear space of dimension (#S1)(\#S-1) given by the vanishing of the coordinates indexed by the complement $\S[r+1]\backslash$ S.
Definition 3.6.
@@
Definition 3.7.
For $S\subset[r+1]$ of cardinality at most r1r-1, let $\mathsf{E}_{S}$ be the class of the strict transform of $\Lambda_{S}$. Let $\mathsf{H}^{\vee},\mathsf{E}_{S}^{\vee}\in H_{2}(Y)$ be the basis of 1-cycles dual to the basis $\mathsf{H},\mathsf{E}_{S}\in H^{2}(Y)$ of divisors.
@@
A point $y\in Y$ can be expressed in coordinates as follows. First, take a point $y_{0}\in\mathbb{P}^{r}=\mathbb{P}(\mathbb{C}[[r+1]]),$ where $\mathbb{C}[[r+1]]$ denotes the vector space with basis given by the set $r+1]=\{1,2,\ldots,r+1\}$. The point $y_{0}$ is the image of $y$ under the composite blow-up $Y\to\mathbb{P}^{r}$. Then, let $S^{1}\subsetneq[r+1]=:S^{0}$ be the subset of coordinates of $y_{0}$ which are equal to zero, corresponding to the minimal $T$-invariant subvariety of $\mathbb{P}^{r}$ in which $y_{0}$ lies. Then, let $y_{1}\in\mathbb{P}(\mathbb{C}[S^{1}])\cong\mathbb{P}^{\#S^{1}-1}$ be a point, representing a projectivized normal vector to $\Lambda_{(S^{1})^{c}}$ at $y_{0}$. Define $S^{2}\subsetneq S^{1}$ analogously, as the set of coordinates of $y_{1}$ equal to zero, and continue until $y_{k}$ has all coordinates non-zero. Then, $y$ consists of the data of the points $(y_{0},y_{1},\ldots,y_{k})$.
@@
Let $C$ be a smooth curve. A map $f:C\to Y$, which, upon post-composition with the blow-up $Y\to\mathbb{P}^{r}$, has image not contained in any torus-invariant subvariety, is given by the following data.
@@
  • for each $S\subset[r+1]$ with $1\leq\#S\leq r-1$, an effective divisor $D_{S}\subset C$, and
@@
The point $f(p)=(y_{0},\ldots,y_{k})$ may be computed as follows. First $y_{0}$ is the point $f_{0}(p):\cdots:f_{r+1}(p)],$ where the $f_{j}$ are regarded as sections of $\mathcal{L}$. Then, the subset $S^{1}\subset[r+1]$ is the set of indices $j$ for which $f_{j}(p)=0$. The point $y_{1}$ has coordinates given by the values of $f_{j}(p)$ after twisting $\mathcal{L}$ down by the unique positive multiple of $p$ for which the $f_{j}(p)$ are well-defined and not all zero for $j\in S^{1}$. The rest of the $y_{2},\ldots,y_{k}$ are then determined by further twists of $\mathcal{L}$.
@@ 3.1.2. Transversality
In this section, we prove Proposition 3.2. We continue to assume $\ell=r+1$. The main difficulty is to show that a point of $V(x_{1},\ldots,x_{n})$ represents an “honest” map in class $\beta$. More precisely, let $\mathbb{P}^{\circ}\subset\mathbb{P}$ be the open locus of $f=[f_{1}:\cdots:f_{r+1}]$ for which the $f_{j}$ share no common zeroes, when simultaneously viewed as sections of $\mathcal{L}$. Then, we wish to show that $V(x_{1},\ldots,x_{n})\subset\mathbb{P}^{\circ}$.
@@
Remark 3.9.
We will later use the assumption $n-d\geq g+1$ again (see Lemma 3.18), but we point out that it is crucial for Lemma 3.8 to hold, which in turn is necessary for the transversality.
@@
On the other hand, when $g=0$ and $r=2$, one can check that the assumption $n-d\geq 1$ can be dropped, also later in Lemma 3.18.
We now turn to the main argument. The strategy is as follows: for any initial $f_{\text{init}}\in V(x_{1},\ldots,x_{n})$, we describe an algorithm that essentially amounts to twisting down at base-points of $f_{\text{init}}$ until we are able to define a map $f:C\to\mathbb{P}^{r}$. The map $f$ will also have various incidence conditions with respect to the torus-invariant loci of $\mathbb{P}^{r}$ and the $p_{i},x_{i}$; a naive dimension count would immediately predict the nonexistence of such a map, unless $f\in\mathbb{P}^{\circ}$ to begin with. We show that this expectation holds by passing from $f$ to a map from $f:C\to Y$, where $Y$ is the permutohedral variety of the previous section, moving in a family of the expected dimension by Lemma 1.1.
For the initial point $f_{\text{init}}\in V(x_{1},\ldots,x_{n})$, denote the underlying line bundle by $\mathcal{L}_{\text{init}}$, the underlying divisors by $D_{j}$, and the underlying sections by
$(f_{j})_{\text{init}}\in H^{0}\left(C,\mathcal{L}_{\text{init}}\left(-\sum_{v\neq j}D_{v}\right)\right).$
Our procedure consists of a sequence of modifications to the following data.
@@
  • For all $j=1,2,\ldots,r+1$, a non-zero (see Lemma 3.8) section
@@
The notation $\mathcal{L},f_{j}$ contrasts with $\mathcal{L}_{\text{init}},(f_{j})_{\text{init}}$, because $\mathcal{L},f_{j}$ will be modified throughout the algorithm, whereas $\mathcal{L}_{\text{init}},(f_{j})_{\text{init}}$ are fixed throughout. Similarly, when $S=\{j\}$, we contrast the divisor $D_{\{j\}}$, which is modified throughout the algorithm, with $D_{j}$, which is fixed. The letter $f$ is only used at the end when we obtain a map $f:C\to\mathbb{P}^{r}$ after no base-points remain.
The data is initialized as follows.
  • $\mathcal{L}\leftarrow\mathcal{L}_{\text{init}}$,
  • $D_{\{j\}}\leftarrow D_{j}$ for $j=1,\ldots,r+1$, and $D_{S}\leftarrow 0$ if #S>1\#S>1, and
  • $f_{j}\leftarrow(f_{j})_{\text{init}}$.
We will repeatedly consider the following conditions on $f_{j}$ given $p\in C$.
@@ 1. Step 1
    Let $\alpha>0$ be the largest order to which $p$ vanishes at every $f_{j}$ (as a section of $\mathcal{L}(-\sum_{v\neq j}D_{\{v\}})$) to order $\alpha$. Then, make the following modifications.
@@
    • each $D_{S}$ stays the same (in particular, all $D_{S}$ with #S>1\#S>1 remain empty), and
@@
      (In other words, $f_{j}$ is “twisted down by $\alpha$ at pp.”)